# Deep Learning for Multivariate Time Series Imputation: A Survey

### v2

@@ Abstract
Missing values are ubiquitous in multivariate time series (MTS) data, posing significant challenges for accurate analysis and downstream applications. In recent years, deep learning-based methods have successfully handled missing data by leveraging complex temporal dependencies and learned data distributions. In this survey, we provide a comprehensive summary of deep learning approaches for multivariate time series imputation (MTSI) tasks. We propose a novel taxonomy that categorizes existing methods based on two key perspectives: imputation uncertainty and neural network architecture. Furthermore, we summarize existing MTSI toolkits with a particular emphasis on the PyPOTS Ecosystem, which provides an integrated and standardized foundation for MTSI research. Finally, we discuss key challenges and future research directions, which give insight for further MTSI research. This survey aims to serve as a valuable resource for researchers and practitioners in the field of time series analysis and missing data imputation tasks.
@@ 1 Introduction
The data collection process of multivariate time series in various fields is often fraught with difficulties and uncertainty. In IoT systems, sensor failures and unstable environments lead to missing measurements Li et al. (2023c). Clinical studies face challenges from irregular sampling and privacy concerns Ibrahim et al. (2012); Esteban et al. (2017); Qian et al. (2024). Financial and transportation systems encounter data gaps due to system downtime and communication issues Bai and Ng (2008); Gong et al. (2021). These missing values can significantly affect the accuracy and reliability of downstream analysis and decision-making. In real-world datasets like PhysioNet2012 Silva et al. (2012b), missing rates can exceed 80%. Consequently, exploring how to reasonably and effectively impute missing components in multivariate time series is attractive and essential.
@@
To the best of our knowledge, this is the first comprehensive and systematic review of deep learning algorithms in the realm of MTSI, aiming to stimulate further research in this field. A corresponding resource that has been continuously updated can be found on our GitHub repository [^1] https://github.com/WenjieDu/Awesome_Imputation. In summary, the contributions of this survey include:
@@ 3.2 RNN-based Models
As a natural way to model sequential data, Recurrent Neural Networks (RNNs) get developed early on the topic of advanced time-series analysis, and imputation is not an exception. GRU-D Che et al. (2018), a variant of GRU, is designed to process time series containing missing values. It is regulated by a temporal decay mechanism, which takes the time-lag matrix $\mathbf{\delta}_{i}$ as input and models the temporal irregularity caused by missing values. Temporal belief memory Kim and Chi (2018), inspired by a biological neural model called the Hodgkin–Huxley model, is proposed to handle missing data by computing a belief of each feature’s last observation with a bidirectional RNN and imputing a missing value based on its according belief. M-RNN Yoon et al. (2019) is an RNN variant that works in a multi-directional way. This model interpolates within data streams with a bidirectional RNN model and imputes across data streams with a fully connected network. BRITS Cao et al. (2018) models incomplete time series with a bidirectional RNN. It takes missing values as variables of the RNN graph and fills in missing data with the hidden states from the RNN. In addition to imputation, BRITS is capable of working on the time series classification task simultaneously. Both M-RNN and BRITS adopt the temporal decay function from GRU-D to capture the informative missingness for performance improvement. Subsequent works, such as Luo et al. (2018, 2019); Liu et al. (2019); Miao et al. (2021), combine RNNs with the GAN structure to output imputation with higher accuracy.
@@ 3.5 Attention-based Models
Since Transformer is proposed in Vaswani et al. (2017), the self-attention mechanism has been widely used to model sequence data including time series Wen et al. (2023). DeepMVI Bansal et al. (2021) integrates transformers with convolutional techniques, tailoring key-query designs to effectively address missing value imputation. For each time series, DeepMVI harnesses attention mechanisms to concurrently distill long-term seasonal, granular local, and cross-dimensional embeddings, which are concatenated to predict the final output. SAITS Du et al. (2023) employs a self-supervised training scheme to deal with missing data, which integrates dual joint learning tasks: a masked imputation task and an observed reconstruction task. This method, featuring two diagonal-masked self-attention blocks and a weighted-combination block, leverages attention weights and missingness indicators to enhance imputation precision. Besides, ImputeFormer Nie et al. (2024) introduces a novel Transformer-based framework that leverages self-attention and temporal context modeling to accurately recover missing values. In addition to the above models, the attention mechanism is also widely adapted to build the denoising network in diffusion models like CSDI Tashiro et al. (2021), MIDM Wang et al. (2023), PriSTI Liu et al. (2023a), Diffusion-TS Yuan and Qiao (2024) and in GNN-based models like Casper Jing et al. (2024) and HSPGNN Liang et al. (2024).
@@ 4.4 Diffusion-based Models
CSDI, introduced in Tashiro et al. (2021), stands out as the pioneering diffusion model specifically designed for MTSI. Different from conventional diffusion models, CSDI adopts a conditioned training approach, where a subset of observed data is utilized as conditional information to facilitate the generation of the remaining segment of observed data. However, the denoising network in CSDI relies on two transformers, exhibiting quadratic complexity concerning the number of variables and the time series length. This design limitation raises concerns about memory constraints, particularly when modeling extensive multivariate time series. In response to this challenge, a subsequent work by Alcaraz and Strodthoff (2023) introduces SSSD, which addresses the quadratic complexity issue by replacing transformers with structured state space models Gu et al. (2022). This modification proves advantageous, especially when handling lengthy multivariate time series, as it mitigates the risk of memory overflow. Another approach CSBI, introduced in  Chen et al. (2023c), improves the efficiency by modeling the diffusion process as a Schrodinger bridge problem, which could be transformed into computation-friendly stochastic differential equations. Also, SADI Dai et al. (2024) is a similarity-aware diffusion model that leverages a self-attention mechanism to capture inter-patient similarities for effective imputation of missing values. MTSCI Zhou et al. (2024) leverages cross-channel correlations and multi-scale temporal dynamics features to effectively recover missing values.
Moreover, the efficacy of diffusion models is notably influenced by the construction and utilization of conditional information. MIDM Wang et al. (2023) proposes to sample noise from a distribution conditional on observed data’s representations in the denoising process, In this way, it can explicitly preserve the intrinsic correlations between observed and missing data. PriSTI Liu et al. (2023a) introduces the spatiotemporal dependencies as conditional information, i.e., provides the denoising network with spatiotemporal attention weights calculated by the conditional feature for spatiotemporal imputation. Besides, FGTI Yang et al. (2024a) is a frequency-guided framework that leverages spectral analysis to effectively capture both global periodic patterns and local temporal dynamics, thereby enhancing missing data recovery.
@@ 5.1 Pre-Trained Foundation Models
PFMs leverage large-scale pretraining on diverse datasets to enhance generalization and adaptability Talukder et al. (2024). For example, by pretraining on a vast collection of multivariate time series, Timer Liu et al. (2024) learns a rich set of temporal representations, enabling it to generalize across different domains. It employs a self-supervised contrastive learning framework that refines its ability to reconstruct missing values while maintaining consistency in long-range dependencies. Building on the idea of large-scale pretraining, MOMENT Goswami et al. (2024) introduces the Time Series Pile and facilitates pretraining of high-capacity Transformer models using masked time series prediction tasks. On the other hand, NuwaTS Cheng et al. repurposes pre-trained language models for time series imputation, utilizing specialized embeddings and contrastive learning to handle various missing data patterns across domains. Meanwhile, Timemixer++ Wang et al. (2024a) explores an alternative to Transformer-based architectures by adopting token-mixing techniques. By mixing temporal and feature-wise representations, Timemixer++ effectively captures both short- and long-term dependencies while reducing computational overhead. These PFMs represent a significant step forward in MTSI, offering scalable, generalizable, and high-performance solutions that adapt to various missing data patterns across domains.
@@ 5.2 Large Language Models
LLMs have shown good capabilities in sequential modeling due to their autoregressive nature, which enables them to capture sequence dependencies through next-token prediction. Their extensive model parameters equip them with a strong learning capacity, making them suitable for handling MTSI tasks. For example, GPT4TS Zhou et al. (2023) optimizes GPT-2 for MTSI by introducing a key architectural adjustment—the freezing of the attention module. This design allows the model to focus on fine-tuning positional embeddings and layer normalization layers using a small number of samples. Such targeted fine-tuning significantly improves the model’s ability to reconstruct missing values with minimal data availability. Furthermore, LLM-TS Integrator Chen et al. (2024a) enhances LLM-based MTSI tasks by integrating statistical and deep learning methods within a hybrid framework. It introduces an adaptive retrieval mechanism to select relevant historical patterns and a self-correction module for iterative refinement. By leveraging retrieval-augmented generation, the model effectively handles irregular and sparse time-series data.

### v3

@@ Abstract
Missing values are ubiquitous in multivariate time series (MTS) data, posing significant challenges for accurate analysis and downstream applications. In recent years, deep learning-based methods have successfully handled missing data by leveraging complex temporal dependencies and learned data distributions. In this survey, we provide a comprehensive summary of deep learning approaches for multivariate time series imputation (MTSI) tasks. We propose a novel taxonomy that categorizes existing methods based on two key perspectives: imputation uncertainty and neural network architecture. Furthermore, we summarize existing MTSI toolkits with a particular emphasis on the PyPOTS Ecosystem, which provides an integrated and standardized foundation for MTSI research. Finally, we discuss key challenges and future research directions, which give insight for further MTSI research. This survey aims to serve as a valuable resource for researchers and practitioners in the field of time series analysis and missing data imputation tasks. A well-maintained MTSI paper and tool list is available at https://github.com/WenjieDu/Awesome_Imputation.
@@ 1 Introduction
The data collection process of multivariate time series in various fields is often fraught with difficulties and uncertainty. In IoT systems, sensor failures and unstable environments lead to missing measurements Li et al. (2023). Clinical studies face challenges from irregular sampling and privacy concerns Ibrahim et al. (2012); Esteban et al. (2017); Qian et al. (2025). Financial and transportation systems encounter data gaps due to system downtime and communication issues Bai and Ng (2008); Gong et al. (2021). These missing values can significantly affect the accuracy and reliability of downstream analysis and decision-making. In real-world datasets like PhysioNet2012 Silva et al. (2012), missing rates can exceed 80%. Consequently, exploring how to reasonably and effectively impute missing components in multivariate time series is attractive and essential.
@@
To the best of our knowledge, this is the first comprehensive and systematic review of deep learning algorithms in the realm of MTSI, aiming to stimulate further research in this field. A corresponding resource that has been continuously updated can be found in our GitHub repository [^1] https://github.com/WenjieDu/Awesome_Imputation. In summary, the contributions of this survey include:
@@ 3.2 RNN-based Models
As a natural way to model sequential data, Recurrent Neural Networks (RNNs) were developed early on the topic of advanced time-series analysis, and imputation is not an exception. GRU-D Che et al. (2018), a variant of GRU, is designed to process time series containing missing values. It is regulated by a temporal decay mechanism, which takes the time-lag matrix $\mathbf{\delta}_{i}$ as input and models the temporal irregularity caused by missing values. Temporal belief memory Kim and Chi (2018), inspired by a biological neural model called the Hodgkin–Huxley model, is proposed to handle missing data by computing a belief of each feature’s last observation with a bidirectional RNN and imputing a missing value based on its corresponding belief. M-RNN Yoon et al. (2019) is an RNN variant that works in a multi-directional way. This model interpolates within data streams with a bidirectional RNN model and imputes across data streams with a fully connected network. BRITS Cao et al. (2018) models incomplete time series with a bidirectional RNN. It takes missing values as variables of the RNN graph and fills in missing data with the hidden states from the RNN. In addition to imputation, BRITS is capable of working on the time series classification task simultaneously. Both M-RNN and BRITS adopt the temporal decay function from GRU-D to capture the informative missingness for performance improvement. Subsequent works, such as Luo et al. (2018, 2019); Liu et al. (2019); Miao et al. (2021), combine RNNs with the GAN structure to output imputation with higher accuracy.
@@ 3.5 Attention-based Models
Since Transformer is proposed in Vaswani et al. (2017), the self-attention mechanism has been widely used to model sequence data including time series Wen et al. (2023). DeepMVI Bansal et al. (2021) integrates transformers with convolutional techniques, tailoring key-query designs to effectively address missing value imputation. For each time series, DeepMVI harnesses attention mechanisms to concurrently distill long-term seasonal, granular local, and cross-dimensional embeddings, which are concatenated to predict the final output. SAITS Du et al. (2023) employs a self-supervised training scheme to deal with missing data, which integrates dual joint learning tasks: a masked imputation task and an observed reconstruction task. This method, featuring two diagonal-masked self-attention blocks and a weighted-combination block, leverages attention weights and missingness indicators to enhance imputation precision. Besides, ImputeFormer Nie et al. (2024) introduces a novel Transformer-based framework that leverages self-attention and temporal context modeling to accurately recover missing values. In addition to the above models, the attention mechanism is also widely adapted to build the denoising network in diffusion models like CSDI Tashiro et al. (2021), MIDM Wang et al. (2023), PriSTI Liu et al. (2023), Diffusion-TS Yuan and Qiao (2024) and in GNN-based models like Casper Jing et al. (2024) and HSPGNN Liang et al. (2024a).
@@ 4.4 Diffusion-based Models
CSDI, introduced in Tashiro et al. (2021), stands out as the pioneering diffusion model specifically designed for MTSI. Different from conventional diffusion models, CSDI adopts a conditioned training approach, where a subset of observed data is utilized as conditional information to facilitate the generation of the remaining segment of observed data. However, the denoising network in CSDI relies on two transformers, exhibiting quadratic complexity concerning the number of variables and the time series length. This design limitation raises concerns about memory constraints, particularly when modeling extensive multivariate time series. In response to this challenge, a subsequent work by Alcaraz and Strodthoff (2023) introduces SSSD, which addresses the quadratic complexity issue by replacing transformers with structured state space models Gu et al. (2022). This modification proves advantageous, especially when handling lengthy multivariate time series, as it mitigates the risk of memory overflow. Another approach CSBI, introduced in  Chen et al. (2023), improves the efficiency by modeling the diffusion process as a Schrodinger bridge problem, which could be transformed into computation-friendly stochastic differential equations. Also, SADI Dai et al. (2024) is a similarity-aware diffusion model that leverages a self-attention mechanism to capture inter-patient similarities for effective imputation of missing values. MTSCI Zhou et al. (2024) leverages cross-channel correlations and multi-scale temporal dynamics features to effectively recover missing values.
Moreover, the efficacy of diffusion models is notably influenced by the construction and utilization of conditional information. MIDM Wang et al. (2023) proposes to sample noise from a distribution conditional on observed data’s representations in the denoising process, In this way, it can explicitly preserve the intrinsic correlations between observed and missing data. PriSTI Liu et al. (2023) introduces the spatiotemporal dependencies as conditional information, i.e., provides the denoising network with spatiotemporal attention weights calculated by the conditional feature for spatiotemporal imputation. Besides, FGTI Yang et al. (2024a) is a frequency-guided framework that leverages spectral analysis to effectively capture both global periodic patterns and local temporal dynamics, thereby enhancing missing data recovery.
@@ 5.1 Pre-Trained Foundation Models
PFMs leverage large-scale pretraining on diverse datasets to enhance generalization and adaptability Liang et al. (2024b). For example, by pretraining on a vast collection of multivariate time series, Timer Liu et al. (2024) learns a rich set of temporal representations, enabling it to generalize across different domains. It employs a self-supervised contrastive learning framework that refines its ability to reconstruct missing values while maintaining consistency in long-range dependencies. Building on the idea of large-scale pretraining, MOMENT Goswami et al. (2024) introduces the Time Series Pile and facilitates pretraining of high-capacity Transformer models using masked time series prediction tasks. On the other hand, NuwaTS Cheng et al. repurposes pre-trained language models for time series imputation, utilizing specialized embeddings and contrastive learning to handle various missing data patterns across domains. Meanwhile, Timemixer++ Wang et al. (2024) explores an alternative to Transformer-based architectures by adopting token-mixing techniques. By mixing temporal and feature-wise representations, Timemixer++ effectively captures both short- and long-term dependencies while reducing computational overhead. These PFMs represent a significant step forward in MTSI, offering scalable, generalizable, and high-performance solutions that adapt to various missing data patterns across domains.
@@ 5.2 Large Language Models
LLMs have shown good capabilities in sequential modeling due to their autoregressive nature, which enables them to capture sequence dependencies through next-token prediction. Their extensive model parameters equip them with a strong learning capacity, making them suitable for handling MTSI tasks Jin et al. (2023). For example, GPT4TS Zhou et al. (2023) optimizes GPT-2 for MTSI by introducing a key architectural adjustment—the freezing of the attention module. This design allows the model to focus on fine-tuning positional embeddings and layer normalization layers using a small number of samples. Such targeted fine-tuning significantly improves the model’s ability to reconstruct missing values with minimal data availability. Furthermore, LLM-TS Integrator Chen et al. (2024) enhances LLM-based MTSI tasks by integrating statistical and deep learning methods within a hybrid framework. It introduces an adaptive retrieval mechanism to select relevant historical patterns and a self-correction module for iterative refinement. By leveraging retrieval-augmented generation, the model effectively handles irregular and sparse time-series data.